OpenStudy (anonymous):

Medal! Simplify. Your answer should contain only positive exponents with no fractional exponents in the denominator.

OpenStudy (anonymous):

\[4m^\frac{ 4 }{ 3 }*m^\frac{ -7 }{ 4 }\]

Nnesha (nnesha):

\[\huge\rm x^m \times x^n = x^{m+n}\] exponent rule! when u multiply same bases you should ad their exponents

OpenStudy (anonymous):

i don't now how to add the fractions. because they've got different denominators

Nnesha (nnesha):

multiply them

Nnesha (nnesha):

\[\frac{ 4 }{ 3 } + \frac{ -7 }{ 4 } = \frac{ ? }{ 3(4) }\]

OpenStudy (anonymous):

\[\frac{ -28 }{ 12 }\]

Nnesha (nnesha):

 4 and 3 are the denominator first fraction doesn't have 4 at the denominator so multiply 4 times 4 2nd fraction -7 doesn't have 3 at the denominator so multiply it by 3 same like cross multiplication but don't forget the signs